Β Verify Stokes’ theorem for 𝐴 = 2π‘₯ βˆ’ 𝑦 𝑖 βˆ’ 𝑦𝑧2 𝑗 βˆ’ 𝑦2π‘§π‘˜, where S is the upper half of the sphere π‘₯2+ 𝑦2 + 𝑧2 = 1 and C is its boundary.


1
Expert's answer
2021-05-27T18:07:45-0400

"\int_C\vec F\cdot \vec{dr}=\int\int_S curl \vec F\cdot \hat ndS\ \ (Stoke's\ Theorem)"


where C is the circle: "x^2+y^2=1,\ z=0\ \ (xy-plane)"


i.e. "x=cost,\ y=sint,\ z=0"

"r=x\vec i+y\vec j\ where\ \ 0\leq \theta\leq 2\pi\\\vec{dr}=dx\vec i+dy\vec j"


where,

"\vec F=(2x-y)\vec i-yz^2\vec j-y^2z\vec k\\\therefore\vec F\cdot \vec{dr}=(2x-y)dx\ (since\ z=0 )"


LHS "= \int_C \vec F\cdot \vec{dr}=\int_C(2x-y)dx"


"=\int_0^{2\pi}(2cost-sint)(-sint)dt\\\ \\=\int_0^{2\pi}(sin^2t-2costsint)dt\\\ \\=\int_0^{2\pi}\{\frac{1}{2}(1-cos2t)-sin2t\}dt\\\ \\=[\dfrac{t}{2}-\dfrac{sin2t}{4}-\dfrac{cos2t}{2}]_o^{2\pi}\\\ \\=(\dfrac{1}{2}-\dfrac{1}{2})+(\pi-0)\\\ \\=\pi"


Hence, "\vec F\cdot \vec{dr}=\pi\ \ \ .....(1)"


Also,

"curl \ \ \vec F=\triangledown \times \vec F"

"= \begin{vmatrix}
 \vec i & \vec j&\vec k \\
 \frac{\delta}{\delta x} & \frac{\delta}{\delta y}&\frac{\delta}{\delta z}\\ 2x-y&-yz^2&-y^2z
\end{vmatrix}\\\ \\=\vec i(-2yz+2yz)-\vec j (0)+\vec k (0+1)\\=\vec k"


"\therefore \vec {dS}=\hat n dS\\=dydz\ i+dzdx\ j+dxdy\ k"


Hence,

"RHS\ =\ \int\int_S\ curl\ \vec F\cdot \hat {n}dS=\int\int dxdy=\pi\ \ .....(2)"


"Since,\ \ \int\int dxdy" represent the area of the circle "x^2+y^2=1" which is "\pi"


Thus, from (1) and (2) we conclude that the theorem is verified.